# Scaling RL to Long Videos

### v1

@@ Abstract
Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning. We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 52K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling. In experiments, LongVILA-R1-7B achieves strong performance on long video QA benchmarks such as VideoMME. It also outperforms Video-R1-7B and even matches Gemini-1.5-Pro across temporal reasoning, goal and purpose reasoning, spatial reasoning, and plot reasoning on our LongVideo-Reason-eval benchmark. Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training. LongVILA-R1 demonstrates consistent performance gains as the number of input video frames scales. LongVILA-R1 marks a firm step towards long video reasoning in VLMs. In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models. On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames / around 256k tokens). Code and models are available at https://github.com/NVlabs/Long-RL.
@@ 1 Introduction
In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long-video understandings. Firstly, we strategically construct a high-quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason. Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 52K high-quality Question-Reasoning-Answer pairs for long videos. We use 18K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 33K samples with an additional 110K video data [2] for reinforcement learning. This two-stage training combines high-quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning. We also manually curate a balanced set of 1K long-video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework. LongVideo-Reason comprises a total of 18,077 videos, from which 52K QAs with reasoning annotations. Additionally, we also include additional 110K QAs from existing works [12, 13, 14, 15, 16] for RL training.
@@
LongVILA-R1-7B shows strong performance on general video benchmark, e.g., 68.4% on VideoMME [18](with subtitle). On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 67.9% across four reasoning categories, surpassing open-source models including Video-R1-7B [2] and proprietary models such as GPT-4o [19] by a large margin, and matches the performance of Gemini-1.5-Pro [20]. In addition, LongVILA-R1 exhibits progressive performance improvements as the number of input frames increased.
@@ 3 LongVideo-Reason Data Construction
High-quality annotated datasets for long videos are critical for VLMs. Existing publicly available long-video datasets lack high-quality reasoning annotations. In this section, we detail the data construction process (Figure 4) of a 52K long-video QAs with reasons.
@@ 3.1 Overview of Data Curation
We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left). We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 52K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle). This dataset is designed to support various types of long-video reasoning tasks comprehensively.
Given the sensitivity of GRPO to batch sampling [41, 25], a data filtering approach is adopted. Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets. Questions consistently answered correctly or incorrectly are labeled as "easy" or "hard" while those inducing diverse predictions are labeled as "medium". We use both the "easy" and "hard" samples for Stage-1 COT-SFT and use the "medium" samples for Stage-2 RL training. The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers. The COT-SFT subset (18K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities. Meanwhile, the RL subset contains 33K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning. To further enhance RL scaling, we incorporate an additional 110K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16]. This combination improves the model’s generalization.
@@ 3.2 Long-Video Reasoning Generation
We introduce an automated annotation pipeline (Figure 4) that generates high-quality Question-Reasoning-Answer pairs from long videos. This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions. Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video. Specifically, we design four types of prompts to encourage the LLM to generate a Question-Reasoning-Answer pair that focuses on one of the four types of reasoning: Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning. To ensure VLMs focus on visual details, we also craft the prompts with phrases such as “checking the video” and “analyzing the scene”, which guide iterative examination of visual content. Finally, an LLM is then used to refine and streamline the reasoning steps. We also manually curated 1,000 high-quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities. This entire data procedure consumes about 40,000 GPU hours on H100s.
@@ 4 LongVILA Training Pipeline
As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 18K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
@@ 4.1 Long Video CoT Supervised Fine-Tuning
Utilizing 52K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 18K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL. This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios. To efficiently perform SFT on hundreds of frames, we adopt the MM-SP [4] training system from LongVILA. As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with foundational reasoning capabilities.
@@ 4.2 GRPO for Long Video
Considering the sensitivity of GRPO to sampling during training [41], we use the 33K filtered data for reinforcement learning as described in Section 3.1. Additionally, an extra 110K samples from Video-R1 [2] are incorporated to scale up the RL. This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
@@ 5 Multi-modal Reinforcement SP
Table 1: Performance on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18]. LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
Model
ActNet-QA
EgoSchema
EventBench
LVideoBench
PercepTest
MVBench
NExT-QA
VNBench
VideoMME
test test val val val test mc val w/o sub. w/ sub.
GPT-4V [49] 57.0 - 32.6 61.3 - 43.5 - - 59.9 63.3
GPT-4o [19] - - 53.3 66.7 - - - 64.4 71.9 77.2
Gemini-1.5-Pro [20] 57.5 72.2 43.2 64.0 - - - 66.7 75.0 81.3
Video-LLaVA-7B [50] 45.3 38.4 5.9 37.6 - 43.5 - 12.4 39.9 41.6
Flash-VStream-7B [51] 51.9 - - - - - 61.6 - - -
ShareGPT4Video-8B [52] 50.8 - - 41.8 - 51.2 - - 39.9 43.6
VideoLLaMA2-7B [53] 50.2 51.7 6.9 - 51.4 54.6 - 4.5 47.9 50.3
VideoLLaMA2.1-7B [53] 53.0 53.1 - - 54.9 57.3 - - 54.9 56.4
Kangaroo-8B [54] - 62.7 - 54.8 - 61.1 - - 56.0 57.6
PLLaVA-7B [55] 56.3 - 28.2 39.2 - 46.6 - - - -
LLaVA-OV-7B [56] 56.7 60.1 - 56.4 57.1 56.7 79.4 51.8 58.2 61.5
LongVILA-7B [4] 59.5 67.7 58.0 57.1 58.1 67.1 80.7 63.0 60.1 65.1
LongVILA-R1-7B 59.6 68.1 58.2 57.6 59.3 67.6 80.8 66.8 62.4 68.4
Table 2: Performance on LongVideo-Reason-eval. LongVILA-R1-7B achieves the highest overall score.
@@
     GPT-4o [19]      64.4      62.4      60.4      55.6      60.7
     Gemini-1.5-Pro [20]      73.2      66.4      66.4      63.2      67.3
     Video-R1-7B [2]      66.4      62.8      63.6      58.0      62.7
     LongVILA-R1-7B      71.6      66.4      63.6      70.0      67.9
@@
LongVILA-R1-7B 62.4 71.9 62.0 53.3 68.4 76.2 69.8 59.1
5.1 Stage 1 - Rollout with Paralleled Encoding
5.2 Stage 2 - Prefilling with Sequence Parallelism
@@ 6.1 Main Results
Table 1 shows the performance comparison on 9 video benchmarks [43, 44, 45, 46, 14, 47, 13, 48, 18]. LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks. Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [50, 57, 52, 58, 53, 59, 54, 60, 1, 2] under comparable model sizes on the Video-MME [18] benchmark. The LongVILA-R1-7B achieves the leading scores under different video lengths. LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 60.3 and 65.9 in the settings without subtitles and with subtitles, respectively. Table 2 compares the results of our LongVideo-Reason-eval benchmark. We also show some qualitative results in the appendix. LongVILA-R1-7B model achieves a strong performance with an average score of 67.9, significantly surpassing Video-R1-7B [2] and GPT-4o [19] and slightly outperforms Gemini-1.5-Pro [20]. While Gemini-1.5-Pro demonstrates strong performance in Temporal Reasoning, Goal and Purpose Reasoning, and Plot and Narrative Reasoning, the LongVILA-R1-7B reaches scores that are generally comparable. Notably, in the Spatial Reasoning category, LongVILA-R1-7B achieves a score of 70.0.
@@ 6.2 Ablation Study
The reasoning capability of LongVILA-R1 scales consistently with the number of input video frames. Specifically, Figure 8 illustrates the performance of LongVILA-1.5B (grey line) and LongVILA-1.5B-R1 (red line) on the long-video reasoning benchmark under varying frame inputs. With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B. However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3. Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process. In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames, due to its inability to effectively process such dense visual information. The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
@@ 7 Conclusion
We present a comprehensive framework designed to fully scale VLMs for reasoning over long videos. LongVILA-R1 encompasses a meticulously constructed large-scale dataset, LongVideo-Reason, and a parallelized training framework, MR-SP. Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL. LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 68.4% on VideoMME with subtitles, and attaining an average accuracy of 67.9% across four reasoning categories in the LongVideo-Reason testing set. LongVILA-R1 maintains consistent performance improvements, extending the feasible frame count from 16 to 512. Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames / 256k tokens) RL training on a single node of 8 A100 GPUs. In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including the VILA and Qwen series), and even image and video generation models. In the appendix, we discuss limitations, border impacts, detailed data generation process, and provide more examples.
Appendix A Limitations
Despite our best efforts to create a high-quality long video reasoning dataset, the definition of reasoning still requires further refinement and a more comprehensive conclusion. Moreover, while our proposed efficient reasoning RL training framework is the first to support video training with hundreds to thousands of frames, real-world scenarios often involve videos with far greater frame counts. Thus, there remains significant potential for improvement in the extraction and learning of ultra-dense visual information.
Appendix B Border Impacts

### v2

@@ Abstract
Abstract: We introduce a full-stack framework that scales up reasoning in vision-language models (VLMs) to long videos, leveraging reinforcement learning. We address the unique challenges of long video reasoning by integrating three critical components: (1) a large-scale dataset, LongVideo-Reason, comprising 104K long video QA pairs with high-quality reasoning annotations across diverse domains such as sports, games, and vlogs; (2) a two-stage training pipeline that extends VLMs with chain-of-thought supervised fine-tuning (CoT-SFT) and reinforcement learning (RL); and (3) a training infrastructure for long video RL, named Multi-modal Reinforcement Sequence Parallelism (MR-SP), which incorporates sequence parallelism and a vLLM-based engine tailored for long video, using cached video embeddings for efficient rollout and prefilling. In our experiments, LongVILA-R1-7B achieves strong performance on video benchmarks, reaching 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively, and consistently outperforming LongVILA-R1 across multiple benchmarks. Moreover, LongVILA-R1 shows steady performance improvements as the number of input video frames increases. Notably, our MR-SP system achieves up to 2.1$\times$ speedup on long video RL training. In addition, we release our training system for public availability that supports RL training on various modalities (video, text, and audio), various models (VILA and Qwen series), and even image and video generation models. On a single A100 node (8 GPUs), it supports RL training on hour-long videos (e.g., 3,600 frames / around 256k tokens). Code and models are available at https://github.com/NVlabs/Long-RL and https://huggingface.co/Efficient-Large-Model/LongVILA-R1-7B.
@@ 1 Introduction
In this work, we introduce LongVILA-R1, a comprehensive framework exploring the reasoning capabilities for long-video understandings. Firstly, we strategically construct a high-quality dataset with CoT annotations for long video reasoning, named LongVideo-Reason. Leveraging a powerful VLM (NVILA-8B) [11] and a leading open-source reasoning LLM, we develop a dataset comprising 104K high-quality Question-Reasoning-Answer pairs for long videos. We use 36K high-quality samples for Long-CoT-SFT to initialize the model’s reasoning and instruction-following abilities, and 68K samples with an additional 102K video data [12, 13, 14, 15, 16] for reinforcement learning. This two-stage training combines high-quality reasoning annotations with reinforcement learning, enabling LongVILA-R1 to achieve superior and generalized video reasoning. We also manually curate a balanced set of 1K long-video samples to build a new benchmark, LongVideo-Reason-eval, that evaluates performance from four perspectives: Temporal, Goal and Purpose, Spatial, and Plot and Narrative, for a comprehensive assessment.
Figure 3: Data Distribution of LongVideo-Reason and total data in the LongVILA-R1 training framework. LongVideo-Reason comprises a total of 18K videos, from which 104K QAs with reasoning annotations. Additionally, we also include additional 102K QAs from existing works [12, 13, 14, 15, 16] for RL training.
@@
In our experiments, LongVILA-R1-7B demonstrates strong performance on video benchmarks, achieving 65.0% and 70.7% accuracy on VideoMME without and with subtitles, respectively. It consistently outperforms LongVILA-R1 across a range of benchmarks, including ActivityNet-QA, LongVideoBench, PerceptionTest, NExT-QA, VNBench, and VideoMME. Additionally, LongVILA-R1 exhibits steady performance gains as the number of input video frames increases from 16 to 512 frames. On our LongVideo-Reason-eval benchmark, LongVILA-R1-7B achieves an average accuracy of 72.0% across four reasoning categories—surpassing open-source models like Video-R1-7B[2], as well as proprietary models such as Gemini-1.5-Pro[18].
@@ 3 LongVideo-Reason Data Construction
High-quality annotated datasets for long videos are critical for VLMs. Existing publicly available long-video datasets lack high-quality reasoning annotations. In this section, we detail the data construction process (Figure 4) of a 104K long-video QAs with reasons.
@@ 3.1 Overview of Data Curation
We first curate 18K long videos from the Shot2Story dataset [40] (Figure 3, left). We then apply a high-quality automated annotation pipeline for CoT as detailed in Section 3.2, and end up with a total of 104K Question-Reasoning-Answer pairs where each sample, based on the type of question it is reasoning about, can be categorized into Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning (Figure 3, middle). This dataset is designed to support various types of long-video reasoning tasks comprehensively.
Given the sensitivity of GRPO to batch sampling [41, 25], a data filtering approach is adopted. Specifically, we employ a test-scaling method where LongVILA [4] performs inference 10 times on the original datasets. Questions consistently answered correctly or incorrectly are labeled as easy or hard while those inducing diverse predictions are labeled as medium. We filter out too easy and too hard questions, and use remaining questions for training. The reason is that GRPO expects different rollouts of each sample to be diverse in order to have meaningful advantages, and the gradient vanishes if all the rollouts predict correct or incorrect answers. The COT-SFT subset (36K) features high-quality CoT reasoning processes formatted in a standard <think></think><answer></answer> structure, providing abundant resources for warm-up training during Stage 1 of the model’s reasoning capabilities. Meanwhile, the RL subset contains 68K challenging long-video Q&A, which is leveraged in Stage 2 for scaling reasoning through reinforcement learning. To further enhance RL scaling, we incorporate an additional 102K high-quality open-source videos (Figure 3, right) from other datasets [12, 13, 14, 15, 16]. This combination improves the model’s generalization.
@@ 3.2 Long-Video Reasoning Generation
We introduce an automated annotation pipeline (Figure 4) that generates high-quality Question-Reasoning-Answer pairs from long videos. This process begins by segmenting videos into short clips (~10 seconds each), each of which is annotated using the NVILA-8B [11] model to provide descriptive captions. Leveraging the breakthrough in text-based reasoning, we then deploy a leading open-source reasoning LLM, provide the captions of all the clips in each video, and prompt it to generate diverse types of Question-Reasoning-Answer pairs that involve reasoning over the content across the whole video.
Specifically, we design four types of prompts to encourage the LLM to generate a Question-Reasoning-Answer pair that focuses on one of the four types of reasoning: Temporal Reasoning, Goal and Purpose Reasoning, Spatial Reasoning, or Plot and Narrative Reasoning. To ensure VLMs focus on visual details, we also craft the prompts with phrases such as “checking the video” and “analyzing the scene”, which guide iterative examination of visual content. Finally, an LLM is then used to refine and streamline the reasoning steps. We also manually curated 1,000 high-quality complex reasoning questions across four reasoning categories to serve as a new benchmark (LongVideo-Reason-eval) for evaluating VLMs in reasoning abilities. Note that we also include open-ended questions. Among the total 104K QA pairs, approximately half are multiple-choice and the other half are open-ended. This entire data procedure consumes about 80,000 GPU hours on H100s.
@@ 4 LongVILA Training Pipeline
As shown in Figure 5, there are two extended training stages in LongVILA-R1, i.e., (1) warm-up for long video reasoning, utilizing 36K data with high-quality CoT for SFT on the MM-SP system [4]; (2) reinforcement learning with dense frames from long videos.
@@ 4.1 Long Video CoT Supervised Fine-Tuning
Utilizing 104K high-quality question-reasoning-answer pairs, we apply the data filtering method described in Section 3.1 to select 36K examples for long CoT-SFT, serving as a warm-up phase for subsequent RL. This stage equips the model with fundamental reasoning abilities and instruction-following skills for long video scenarios. To efficiently perform SFT on hundreds of frames, we adopt the MM-SP [4] training system from LongVILA. As demonstrated in Section 6.2, SFT solely on our LongVideo-Reason dataset also effectively improves the model with foundational reasoning capabilities.
@@ 4.2 GRPO for Long Video
Considering the sensitivity of GRPO to sampling during training [41], we use the 68K filtered data for reinforcement learning as described in Section 3.1. Additionally, an extra 102K samples from other datasets [12, 13, 14, 15, 16] are incorporated to scale up the RL. This approach aims to guide the model in freely exploring and developing more effective and generalized reasoning strategies.
@@ 5 Multi-modal Reinforcement SP
5.1 Stage 1 - Rollout with Paralleled Encoding
5.2 Stage 2 - Prefilling with Sequence Parallelism
Table 1: Performance on 6 video benchmarks [43, 44, 14, 13, 45, 20]. LongVILA-R1-7B outperforms LongVILA-7B across all benchmarks, with varying margins that reflect different levels of reasoning difficulty.
Model ActivityNet-QA LongVideoBench PerceptionTest NExT-QA VNBench VideoMME
test val val mc val w/o sub. w/ sub.
GPT-4V 57.0 61.3 - - - 59.9 63.3
GPT-4o - 66.7 - - 64.4 71.9 77.2
Gemini-1.5-Pro 57.5 64.0 - - 66.7 75.0 81.3
Video-LLaVA-7B 45.3 37.6 - - 12.4 39.9 41.6
Flash-VStream-7B 51.9 - - 61.6 - - -
ShareGPT4Video-8B 50.8 41.8 - - - 39.9 43.6
VideoLLaMA2-7B 50.2 - 51.4 - 4.5 47.9 50.3
VideoLLaMA2.1-7B 53.0 - 54.9 - - 54.9 56.4
Kangaroo-8B - 54.8 - - - 56.0 57.6
PLLaVA-7B 56.3 39.2 - - - - -
LLaVA-OV-7B 56.7 56.4 57.1 79.4 51.8 58.2 61.5
LongVILA-7B 59.5 57.1 58.1 80.7 63.0 60.1 65.1
LongVILA-R1-7B 64.8 58.0 68.9 81.5 75.5 65.0 70.7
Table 2: Performance on LongVideo-Reason-eval set. LongVILA-R1-7B achieves the highest overall score.
@@ 5.2 Stage 2 - Prefilling with Sequence Parallelism
     Video-R1-7B [2]      61.4      85.0      62.0      58.5      68.1
     Gemini-1.5-Pro [18]      65.4      81.9      67.8      53.3      69.3
     LongVILA-R1-7B      68.1      85.7      70.6      53.3      72.0
@@
LongVILA-R1-7B 65.0 76.6 63.6 54.9 70.7 79.0 69.1 64.0
@@ 6.1 Main Results
Table 1 shows the performance comparison on 6 video benchmarks [43, 44, 14, 13, 45, 20]. LongVILA-R1-7B consistently outperforms LongVILA-7B across all benchmarks, with performance gaps that vary according to the complexity of the reasoning tasks. Table 3 presents the general performance of LongVILA-R1, comparing to existing advanced models [46, 47, 48, 49, 50, 51, 52, 53, 1, 2] under comparable model sizes on the Video-MME [20] benchmark. The LongVILA-R1-7B is tested using 512 video frames as inputs. LongVILA-R1-7B achieves leading scores across different video lengths, obtaining scores of 65.0% and 70.7% in the settings without subtitles and with subtitles, respectively. Table 2 compares the results of our LongVideo-Reason-eval benchmark. We also show some qualitative results in the appendix. LongVILA-R1-7B model achieves a strong performance with an average score of 72.0%, significantly surpassing Video-R1-7B [2] and slightly outperforms Gemini-1.5-Pro [18].
@@ 6.2 Ablation Study
The reasoning capability of LongVILA-R1 scales consistently with the number of input video frames. Specifically, Figure 8 illustrates the performance of LongVILA-1.5B (grey line) and LongVILA-1.5B-R1 (red line) on the long-video reasoning benchmark under varying frame inputs. With only 16 input frames, LongVILA-R1-1.5B is close to LongVILA-1.5B. However, as the number of frames increases to 512, LongVILA-R1-1.5B consistently outperforms and eventually achieves a score of 64.3%. Notably, LongVILA-R1-1.5B demonstrates steady performance improvements throughout the scaling process. In contrast, LongVILA-1.5B hits a performance bottleneck with 256 frames, and got a degradation on 512 frames, due to its inability to effectively process such dense visual information. The enhanced reasoning capabilities of LongVILA-R1-1.5B allow it to effectively integrate and infer information from long videos.
@@ 7 Conclusion
We present a comprehensive framework designed to fully scale VLMs for reasoning over long videos. LongVILA-R1 encompasses a meticulously constructed large-scale dataset, LongVideo-Reason, and a parallelized training framework, MR-SP. Leveraging our curated dataset of 52K long video question-reasoning-answer pairs, combined with other open-source video datasets, we adopt a two-stage training process that integrates CoT-SFT and RL. LongVILA-R1-7B demonstrates outstanding performance on mainstream video benchmarks, achieving 65.0% and 70.7% on VideoMME without and with subtitles. LongVILA-R1 maintains consistent performance improvements, extending the feasible frame count from 16 to 512. Notably, our MR-SP leads to a speed up of 2.1$\times$ for long video RL training, supporting hour-level (3,600 frames / 256k tokens) RL training on a single node of 8 A100 GPUs. In addition, we release our training system to the public, which supports RL training across multiple modalities (video, text, and audio), various models (including the VILA and Qwen series), and even image and video generation models. In the appendix, we discuss RL training curves, border impacts, detailed data generation process, and provide more examples.
Limitations
While our RL training system is capable of handling long video frames efficiently (3,600 frames on a single node of A100 GPUs), scaling to significantly longer sequences, more modalities (like include audio for omni VLMs) or large batch sizes would require distributed training across multiple GPU machines. This requires more GPU machines, making it less feasible to run on limited hardware resources.
Appendix A Border Impacts